# Machine learning applications in archaeological practices: a review

### v3

@@ 4.2.6 Taphonomic classification
When the data and code is made available, such as in Abellán et al. [2022], concerns are easier to both discard and confirm. For instance, the Ensemble.ipynb file provided in the Supplementary Data (henceforth SD) 2 contains an off-by-one error, which has the effect of excluding the last variable when training the models used. If this is indeed the code used for the main article (it is at least for the Supplementary Text 4), it would have considerable implications for the reported results and the accuracy of the reported methods. Moreover, the R file croc paper 2022 code.r contains additional problems for the “correct” bootstrapping section (module 4 in the file), and we could not run the code for this module without errors (with R 4.3.2, and caret 7.0-1).
@@
Other issues may require even more technical knowledge to identify, such as those in Moclán et al. [2019], who evaluated their models with both a bootstrapped and non-bootstrapped experimental sample, seeking to predict the agent (anthropic direct percussion, spotted hyena (C. crocuta) carnivory and wolf (C. lupus) carnivory) of bone fracture planes in bone fragments (Moclán et al. 2019). However, the dataset provided in the Electronic supplementary material 1 (ESM 1) also contains information on the bone fragment the fracture planes come from (e.g. epiphysis presence, length, and number of planes), rather than only the information of the plane itself (Moclán et al., 2019, app. ESM 1). In some cases a single bone fragment could have six fracture planes in the dataset, so when splitting the dataset, some input data may have come from the same bone as some testing data, and thus have this shared bone metadata, likely leading to overfitting as all bone fracture planes from the same bone fragment were caused by the same agent [see (Varoquaux and Cheplygina, 2022)]. Though no code was made available, the authors do state that these metadata variables were indeed used (Moclán et al., 2019, p. 4668). In addition, the number of entries in ESM 1 do not match the number of fragments reported for any of the three classes. Even when counting only unique bone fragments instead of planes using ESM 1, the numbers still do not match for any class. Furthermore, for the hyena class, it is also clear some fragments did not have all their fracture planes added to the ESM 1 dataset, since there are an odd number of entries that state the planes per bone fragment is 2, thus implying there is at least one plane that was not included (Moclán et al., 2019).
@@ 4.3 Recommendations and good practices
Data availability and open access to data is often discussed as a big challenge in archaeological research. In general, the entire workflow, from original data collection to publication, should be accessible. Following the FAIR principles (Wilkinson et al., 2016), data can be stored in open access data articles (e.g. Journal of Open Archaeology Data) or in institutional and international platforms (e.g. Zenodo, OSF, FigShare). The code should also be accessible either on an online platform (e.g.GitHub, GitLab) or in the supplementary material. Moreover, all the results, including those not discussed, should be available to the reader. Adopting a FAIR workflow and open-access data gives the opportunity to share methodologies and outcomes to a wider public, and potentially aid in the preservation of cultural heritage (Fisher et al., 2021). The Peer Community Journal Archaeology (PCI Archaeology) provides a good example of a fully accessible and FAIR process, requesting the open publication of all data, scripts, and code used for the published study, as well as a clear and thorough reporting of all the methods used, enough to be independently reproducible. A list of software and platforms adapted for increasing reproducibility in archaeological research, although slightly outdated, is given in supplementary tables of Strupler and Wilkison [2017].
@@
Furthermore, the validity and (especially inter-rater) reliability of the datasets used are in many cases not questioned enough (Tennie, 2023), while it is an essential element for reproducibility. Another issue is noise, the outcomes may be valid and reliable if analysed together, but may have low precision. All these problems will become more urgent once a larger transition to machine learning-based methods has happened and no traditional benchmarks exist to compare the results obtained from the models to a broader archaeological context. Besides the fact that the traditional benchmarks are often poor in validity and reliability themselves, the opacity (cf. black box) of many machine learning model families (especially artificial neural networks, which have recently surged in popularity) renders the whole problem even more difficult to solve. One key solution could reside in the development of explainable artificial intelligence (XAI), which intends to bring more transparency throughout the entire automation process (Barredo Arrieta et al., 2020).
@@
Interdisciplinarity and collaboration pose another challenge. Archaeologists should have the primary say in formulating research questions, even when working with computer scientists (Orton, 1980, p. 15). For example, the archaeology of early modern buildings integrates archaeologists and architects, yet the research questions ought to remain archaeological in nature for an archaeological study. Methodological questions can be developed by computer scientists, but the primary goals must stay aligned with archaeological objectives [Fig. 10].
@@
Another important theoretical issue is the reflection throughout the automation process, which also corresponds to the concept of “theory in, theory out”. As mentioned earlier, the use of the Knowledge Discovery Process (Hörr, 2011; Hörr et al., 2014) represents an innovative and holistic approach to fully understand an archaeological artefact. It involves an initial clustering or unsupervised approach, followed by a supervised approach, with an optional semi-supervised step in between. This process encourages deeper reflection, a more subjective approach and more informed decision making. It is also a matter of taking the time to label the various data, as Klassen et al. state: “Given the nature of archaeological data, it is often difficult or expensive to get “labels,” for things like artifact typologies and site chronologies(Klassen et al., 2018). Therefore, standardised ontologies and semantic consistency are needed to assure the interoperability of the methods and results (Huggett, 2017; Davis, 2020b).
@@
Finally, while researchers can have doubts about the validity of results obtained through some applications of machine learning, as well as the validity of its underlying reasoning and the validity of the way these methods were applied [see Varoquaux and Cheplygina 2022], and all these aspects need to be critically examined, they also have great potential. They can also reveal unseen relationships within a dataset, moving beyond human interpretation to a more pragmatic system. As Ramazzotti [2020, p. 174] states, machine learning in archaeology can recreate “a possible world of other associations of meaning devoid of sources and dispersed information, it exhibits the nuances and complex interrelations and, furthermore, it helps the interpreter codify other associations that were unforeseen (or hidden)”. Some might even consider highly mechanised and automated processes (machine learning being only one of many) as a new step into human technical gesture (Leroi-Gourhan, 2022, p. 74).
@@ 4.4 Trends and perspectives
Firstly, regarding the general development of machine learning in archaeology, the academic background of new students will evolve. We identified four programs that focused on computer science applications in archaeology. The “Computational Archaeology: GIS, Data Science and Complexity” master at UCL in London, “Digital Archaeology” master at York University, the “Digital and Computational Archaeology” formation at Universität zu Köln in Cologne and the “Master in Digital Heritage and Landscape Archaeology” at the University of Cyprus in Nicosia. This number will undoubtedly grow, and machine learning will become as familiar to the next generation of archaeologists as GIS and remote sensing (Agapiou and Lysandrou, 2015) or open source principles (Batist and Roe, 2024) are today. Although not explored in this review, explainable artificial intelligence is likely to be trialled for many applications, as it also enables the archaeologist to see the hidden relationships between their data, processing workflow, and results. Archaeological development of such tool can allow more interpretability either during the process (Labba et al., 2023) or post hoc (Tenzer et al., 2024).
@@
Regarding automatic structure detection, it seems that, without a doubt, these already well represented tasks in our corpus, will likely find economic opportunities. Private and public companies seek to manage costs and risks associated with new installations and construction, and automatic structure detection can help avoid unnecessary excavations. A private company in Ireland has already tested the ADAF model (Čož et al., 2024) to reduce archaeological intervention time during track construction, by predicting unseen and unreferenced archaeological mounds. Archaeological predictive models might also be regarded as potential tools for construction companies or even public administration to regulate their cultural heritage conservation policy. However, it is crucial to use these tools responsibly, acknowledging that machine learning models are not infallible, nor can they replace manual fieldwork. These tools, if applied commercially, ought to be seen as an avenue to minimise or at the very least better predict the impact of archaeological material during construction, but never as an avenue to avoid archaeological intervention. In fact, it might be possible to observe or even prevent the destruction of archaeological material (El-Hajj, 2021), especially with widely-verified and trusted methodologies. Both software developers and archaeologists must clearly define thresholds for feature classification or prediction and provide diverse datasets to ensure comprehensive predictions to avoid overlooking archaeological features.
@@
The future development of new algorithms for classification of artefacts is another key area. Unlike automatic structure detection, where MR-CNNs dominate, there are no standard models in artefact classification. Although ANNs have taken the lead, as it has in other disciplines (Zheng et al., 2021; Osco et al., 2021; Thai, 2022), a few dominant models, whether within the ANN family of methods or not, will likely emerge. ResNet is already very popular and well-suited for image recognition problems (Gualandi et al., 2021; Anichini et al., 2021; Kowlessar et al., 2021; Berganzo-Besga et al., 2021; Jalandoni et al., 2022), despite its relative antiquity (He et al., 2016). Furthermore, the rising amount of data used will favour deep learning models compared to other types of models (e.g. ensemble learning, Bayesian classifier), due to its higher number of parameters (Sarker, 2021). In the case of APMs, ensemble learning methods like tuned RF and MaxEnt, despite being several decades old, are well-suited to APM (MaxEnt in particular), which does not require data points for absence of features (Yaworsky et al., 2020, 2024a, 2024b). These methods will likely remain in use for some years due to the special situation of APMs. An alternative is to employ deep learning models coupled with interpretability tools such as LIME (Ribeiro et al., 2016) or SHAP (Lundberg and Lee, 2017), which were not applied in our corpus, nor (to our knowledge) have yet to be applied in archaeology. One likely solution will be the development of studies combining "traditional" statistical approaches with supervised or unsupervised machine learning approaches to hopefully obtain more readable and interpretable results (Li et al., 2024).
@@
Finally, we have not explored large language models (LLMs) in detail. The release of GPT-3.5 in 2022 opened many new possibilities (Brown et al., 2020). Our review did not cover these developments as we excluded any articles published after 2022. However, LLMs have already shown promising results in archaeology for teaching (Cobb, 2023) and for literature scraping, as well as classification (Agapiou and Lysandrou, 2023; Bellat et al., 2024). LLMs have even been applied to the automatic classification of artefacts with partially successful results (Lapp and Lapp, 2024). In the future, LLMs could also be used for code generation or translation between programming languages. While Unnatural-Code-LLaMA-34B excels in generating new code, GPT-4 is the better option for code translation (Zheng et al., 2024). LLMs could make coding more accessible to a wider audience of archaeologists and are likely to increase the number of publications on machine learning applications in archaeology. They are not without their concerns, however, as many academic institutions (of Tübingen, 2023; University, 2024; of Technology, 2024; Wang et al., 2024) or even European deciders (Commission et al., 2024) have already set up guidelines to control and regulate the use of such models.
@@ 5 Conclusions
Our extensive review of applications of machine learning in archaeological research, which includes both quantitative and qualitative observations, highlighted several key phenomena. Expanding on Bickler’s assertion that: “Machine learning arrives in archaeology(Bickler, 2021), our findings suggest that machine learning has been widely present in archaeology since at least 2019. The adoption of machine learning techniques has been slower than in other fields, likely due to the inherent inertia of the discipline. To understand the diversity of machine learning used in archaeology, we focused primarily on four parameters: archaeological subfields, family of algorithms applied, model evaluation goal, and scientific task. The initial wave of machine learning applications focused on artefact classification and bioarchaeological problems (Nguifo et al., 1997; Toler-Franklin et al., 2010; Ionescu, 2015; Mircea et al., 2015b). Automatic structure detection has strongly developed after 2019, likely due to the proliferation of image recognition algorithms based on convolutional neural networks (CNNs) and segmentation models such as Multi-Scale Region-Based CNNs, especially since 2021. The heterogeneity of models used in archaeology contrasts with other fields such as medical imaging applications, where CNNs and generative adversarial networks (GANs) are predominantly used (Barragán-Montero et al., 2021).
@@
Our work also highlighted some methodological concerns likely arising from the limited background of many archaeologists in machine learning, as well as vice versa for some programmers in relation to archaeological problems. Davis and Douglass have already pointed out this disconnect in remote sensing applications in archaeology, noting a “disconnect between remote sensing applications and anthropological theory(Davis and Douglass, 2020). Another critical aspect is the reliability and variability of the benchmarks themselves, which are often poor, leading to additional problems in evaluating the models (Tennie, 2023). Furthermore, the suitability of the chosen machine learning methodology to answering the scientific question sought after is crucial: in many cases, statistical methods are sufficient to provide an answer, possibly rendering machine learning methods superfluous. To address the challenges of applying these methods to archaeological questions, we have designed a flowchart of suggested best practices [Fig. 10] to help archaeologists develop a coherent and effective approach. This small tool aims to bridge the gap between archaeological research and machine learning, promoting a more integrated and informed application of these technologies in the field.
@@
Regarding the future of machine learning in archaeology, three key phenomena emerge. Firstly, archaeologists’ ongoing training in digital methods will lead to a more comprehensive workflow and new applications in sub-fields of archaeology little represented in our current corpus (e.g., ethnoarchaeology, archaeobotany). Improving skills and knowledge will also reduce possible methodological issues, as provide a more direct interaction between the original research questions and the available avenues to solve them. Secondly, the constant development of new models will leave space for new proof-of-concept studies with more complex and wider datasets, but also allow for results that should improvement in results in some domains (cultural heritage, text analysis, automatic structure detection, and artefact classification). Thirdly, the recent boom in large language models and generative AI will change research in archaeology at different levels, as it does already for other fields of academia (Andersen et al., 2025). Either by allowing the creation of helpful tools for data curation, teaching tools, or multivariate analysis, these models are easy to use and have already been adopted by many in their daily life (though ethical issues remain; see Hagendorff 2024). One trend whose path remains uncertain is the possible formation of a reflexive community for AI and machine learning applications in archaeology (e.g. such as the COST action MAiA: MAIA 2025; or CAA machine learning special interest group), leading to more standardised practices and protocols with ethical guidelines.
@@
Despite the many concerns on applying machine learning in archaeology, these methods can be very powerful tools, able to massively reduce the time, cost, and (whether a positive or not) labour to process, analyse, and predict even large-scale and complex data [e.g. Kochkov et al. 2024]. The existing reported successes of machine learning in other areas of science, and its already increasing number of applications in archaeology in recent years, all show it is cementing its place in the future of the field. We are likely at – if not already in a post-digital area (Gattiglia, 2025) – a turning point in archaeology, as well as in other disciplines, where a new approach – machine learning – has the possibility to permeate all levels of archaeological practice. Like any other statistical tool, machine learning itself is neither inherently good nor bad; its impact on the field of archaeology depends on how it is used (Huggett, 2017). We should be mindful of the thinking derived from post-processual archaeology in the late 1970s (Hodder, 1986), developed in response to processual archaeology (Phillips and Willey, 1953). The answer of this post-processual archaeology was to: “explain the past rather than describing it(Yoffee and Fowles, 2010). Following this perspective, machine learning applications in archaeology should never be limited to data processing or analysis but should be integrated into a broader archaeological reflection with precise questions, albeit knowingly and carefully, making sure its limitations are well known, well reported, and well addressed.
@@
Author contributions: M.B., J.D.O.F., C.T. and T.S. conceived the idea for the manuscript and initiated and designed the research. M.B. and J.D.O.F., collected data and selected the articles. M.B. and J.D.O.F., performed the statistical analysis and interpreted the results. M.B., J.D.O.F., C.T. and T.S. wrote the manuscript, with critical input from all co-authors. M.B., J.D.O.F., J.S.R, R.T-M., C.T. and T.S. reviewed the manuscript.
@@
Acknowledgment: We would like to thank J. Padarian for his help on the scraping strategy.
@@
Funding: This work has received funding from the Deutsche Forshungemeischaft (DFG) Collaborative Research Center (CRC) 1070 “ResourceCultures”, grant agreement n°215859406 and from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program under grant agreement n°714658; (STONECULT project).

### v5

@@ 4.2.6 Taphonomic classification
When the data and code are made available, such as in Abellán et al. [2022], concerns are easier to both discard and confirm. For instance, the Ensemble.ipynb file provided in the Supplementary Data (henceforth SD) 2 contains an off-by-one error, which has the effect of excluding the last variable when training the models used. If this is indeed the code used for the main article (it is at least for the Supplementary Text 4), it would have considerable implications for the reported results and the accuracy of the reported methods. Moreover, the R file croc paper 2022 code.r contains additional problems for the “correct” bootstrapping section (module 4 in the file), and we could not run the code for this module without errors (with R 4.3.2, and caret 7.0-1).
@@
Other issues may require even more technical knowledge to identify, such as those in Moclán et al. [2019], who evaluated their models with both a bootstrapped and non-bootstrapped experimental sample, seeking to predict the agent (anthropic direct percussion, spotted hyena (C. crocuta) carnivory and wolf (C. lupus) carnivory) of bone fracture planes in bone fragments (Moclán et al. 2019). However, the dataset provided in the Electronic supplementary material 1 (ESM 1) also contains information on the bone fragment the fracture planes come from (e.g. epiphysis presence, length, and number of planes), rather than only the information of the plane itself (Moclán et al., 2019, app. ESM 1). In some cases a single bone fragment could have six fracture planes in the dataset, so when splitting the dataset, some input data may have come from the same bone as some testing data, and thus have this shared bone metadata, likely leading to overfitting as all bone fracture planes from the same bone fragment were caused by the same agent [see (Varoquaux and Cheplygina, 2022)]. Though no code was made available, the authors do state that these metadata variables were indeed used (Moclán et al., 2019, p. 4668). In addition, the number of entries in ESM 1 do not match the number of fragments reported for any of the three classes. Even when counting only unique bone fragments instead of planes using ESM 1, the numbers still do not match for any class. Furthermore, for the hyena class, it is also clear that some fragments did not have all their fracture planes added to the ESM 1 dataset, since there are an odd number of entries that state the planes per bone fragment is 2, thus implying there is at least one plane that was not included (Moclán et al., 2019).
@@ 4.3 Recommendations and good practices
Data availability and open access to data are often discussed as big challenges in archaeological research. In general, the entire workflow, from original data collection to publication, should be accessible. Following the FAIR principles (Wilkinson et al., 2016), data can be stored in open access data articles (e.g. Journal of Open Archaeology Data) or in institutional and international platforms (e.g. Zenodo, OSF, FigShare). The code should also be accessible either on an online platform (e.g.GitHub, GitLab) or as supplementary materials. Moreover, all the results, including those not discussed, should be available to the reader. Adopting a FAIR workflow and open-access data gives the opportunity to share methodologies and outcomes to a wider public, and potentially aid in the preservation of cultural heritage (Fisher et al., 2021). The Peer Community Journal Archaeology (PCI Archaeology) provides a good example of a fully accessible and FAIR process, requesting the open publication of all data, scripts, and code used for the published study, as well as a clear and thorough reporting of all the methods used, enough to be independently reproducible. A list of software and platforms adapted for increasing reproducibility in archaeological research, although slightly outdated, is given in supplementary tables of Strupler and Wilkinson [2017].
@@
Furthermore, the validity and (especially inter-rater) reliability of the datasets used are in many cases not questioned enough (Tennie, 2023), while it is an essential element for reproducibility. Another issue is noise, the outcomes may be valid and reliable if analysed together, but may have low precision. All these problems will become more urgent once a larger transition to machine learning-based methods has happened and no traditional benchmarks exist to compare the results obtained from the models to a broader archaeological context. Besides the fact that the traditional benchmarks are often poor in validity and reliability themselves, the opacity (cf. black box in part 4.1) of many machine learning model families (especially artificial neural networks, which have recently surged in popularity) renders the whole problem even more difficult to solve. One key solution could reside in the development of explainable artificial intelligence (XAI), which intends to bring more transparency throughout the entire automation process (Barredo Arrieta et al., 2020).
@@
Interdisciplinarity and collaboration pose another challenge. Archaeologists should have the primary say in formulating research questions, even when working with computer scientists (Orton, 1980, p. 15). For example, the archaeology of early modern buildings integrates archaeologists and architects, yet the research questions ought to remain archaeological in nature for an archaeological study. Methodological questions can be developed by computer scientists, but the primary goals must stay aligned with archaeological objectives [Fig.˜11].
@@
Another important theoretical issue is the reflection throughout the automation process, which also corresponds to the concept of “theory in, theory out”. As mentioned earlier, the use of the Knowledge Discovery Process (Hörr, 2011; Hörr et al., 2014) represents an innovative and holistic approach to fully understand an archaeological artefact. It involves an initial clustering or unsupervised approach, followed by a supervised approach, with an optional semi-supervised step in between. This process encourages deeper reflection, a more subjective approach and more informed decision making. It is also a matter of taking the time to label the various data, as Klassen et al. state: “Given the nature of archaeological data, it is often difficult or expensive to get “labels” for things like artifact typologies and site chronologies(Klassen et al., 2018). Therefore, standardised ontologies and semantic consistency are needed to assure the interoperability of the methods and results (Huggett, 2017; Davis, 2020b).
@@
Finally, while researchers can have doubts about the validity of results obtained through some applications of machine learning, as well as the validity of its underlying reasoning and the validity of the way these methods were applied [see Varoquaux and Cheplygina 2022], and all these aspects need to be critically examined, they also have great potential. They can reveal unseen relationships within a dataset, moving beyond human interpretation to a more pragmatic system. As Ramazzotti [2020, p. 174] states, machine learning in archaeology can recreate “a possible world of other associations of meaning devoid of sources and dispersed information, it exhibits the nuances and complex interrelations and, furthermore, it helps the interpreter codify other associations that were unforeseen (or hidden)”. Some might even consider highly mechanised and automated processes (machine learning being only one of many) as a new step into human technical gesture (Leroi-Gourhan, 2022, p. 74).
@@ 4.4 Trends and perspectives
Firstly, regarding the general development of machine learning in archaeology, the academic background of new students will evolve. We identified four programs that focused on computer science applications in archaeology. The “Computational Archaeology: GIS, Data Science and Complexity” master at UCL in London, the “Digital Archaeology” master at York University, the “Digital and Computational Archaeology” programme at Universität zu Köln in Cologne and the “Master in Digital Heritage and Landscape Archaeology” at the University of Cyprus in Nicosia. This number will undoubtedly grow, and machine learning will become as familiar to the next generation of archaeologists as GIS and remote sensing (Agapiou and Lysandrou, 2015) or open source principles (Batist and Roe, 2024) are today. Although not explored in this review, explainable artificial intelligence is likely to be trialled for many applications, as it also enables the archaeologist to see the hidden relationships between their data, processing workflow, and results. Archaeological development of such tools can allow more interpretability either during the process (Labba et al., 2023) or post hoc (Tenzer et al., 2024).
@@
Regarding automatic structure detection, it seems that, without a doubt, economic opportunities will emerge for such applications. Private and public companies seek to manage costs and risks associated with new installations and construction, and automatic structure detection can help avoid unnecessary excavations. A private company in Ireland has already tested the ADAF model (Čož et al., 2024) to reduce archaeological intervention time during track construction, by predicting unseen and unreferenced archaeological mounds. Archaeological predictive models might also be regarded as potential tools for construction companies or even public administration to regulate their cultural heritage conservation policy. However, it is crucial to use these tools responsibly, acknowledging that machine learning models are not infallible, nor can they replace manual fieldwork. These tools, if applied commercially, ought to be seen as an avenue to minimise or at the very least better predict the impact of archaeological material during construction, but never as a way to avoid archaeological intervention. In fact, it might be possible to observe or even prevent the destruction of archaeological material (El-Hajj, 2021), especially with widely-verified and trusted methodologies. Both software developers and archaeologists must clearly define thresholds for feature classification or prediction and provide diverse datasets to ensure comprehensive predictions to avoid overlooking archaeological features.
@@
The future development of new algorithms for classification of artefacts is another key area. Unlike automatic structure detection, where MR-CNNs dominate, there are no standard models in artefact classification. Although ANNs have taken the lead, as they have in other disciplines (Zheng et al., 2021; Osco et al., 2021; Thai, 2022), a few dominant models, whether within the ANN family of methods or not, will likely emerge. ResNet is already very popular and well-suited for image recognition problems (Gualandi et al., 2021; Anichini et al., 2021; Kowlessar et al., 2021; Berganzo-Besga et al., 2021; Jalandoni et al., 2022), despite its relative antiquity (He et al., 2016). Furthermore, the rising amount of data used will favour deep learning models compared to other types of models (e.g. ensemble learning, Bayesian classifier), due to its higher number of parameters (Sarker, 2021).
In the case of APMs, ensemble learning methods like tuned RF and MaxEnt, despite being several decades old, are well-suited to APM (MaxEnt in particular), which does not require data points for absence of features (Yaworsky et al., 2020, 2024a, 2024b). These methods will likely remain in use for some years due to the special situation of APMs. An alternative is to employ deep learning models coupled with interpretability tools such as LIME (Ribeiro et al., 2016) or SHAP (Lundberg and Lee, 2017), which were not applied in our corpus, and (to our knowledge) have yet to be applied in archaeology. One likely solution will be the development of studies combining "traditional" statistical approaches with supervised or unsupervised machine learning approaches to hopefully obtain more readable and interpretable results (Li et al., 2024).
@@
Finally, we have not explored large language models (LLMs) in detail. The release of GPT-3.5 in 2022 opened many new possibilities (Brown et al., 2020). Our review did not cover these developments as we excluded any articles published after 2022. However, LLMs have already shown promising results in archaeology for teaching (Cobb, 2023) and literature scraping, as well as classification (Agapiou and Lysandrou, 2023; Bellat et al., 2024). LLMs have even been applied to the automatic classification of artefacts with partially successful results (Lapp and Lapp, 2024). In the future, LLMs could also be used for code generation or translation between programming languages. While Unnatural-Code-LLaMA-34B excels in generating new code, GPT-4 is the better option for code translation (Zheng et al., 2024). LLMs could make coding more accessible to a wider audience of archaeologists and are likely to increase the number of publications on machine learning applications in archaeology. They are not without their concerns, however, as many academic institutions (of Tübingen, 2023; University, 2024; of Technology, 2024; Wang et al., 2024) or even European deciders (Commission et al., 2024) have already set up guidelines to control and regulate the use of such models.
@@ 5 Conclusions
Our extensive review of applications of machine learning in archaeological research, which includes both quantitative and qualitative observations, highlighted several key phenomena. Expanding on Bickler’s assertion that: “Machine learning arrives in archaeology(Bickler, 2021), our findings suggest that machine learning has been widely present in archaeology since at least 2019. The adoption of machine learning techniques has been slower than in other fields, likely due to the inherent inertia of the discipline. To understand the diversity of machine learning approaches used in archaeology, we focused primarily on four parameters: archaeological subfields, family of algorithms applied, model evaluation goal, and scientific task. The initial wave of machine learning applications focused on artefact classification and bioarchaeological problems (Nguifo et al., 1997; Toler-Franklin et al., 2010; Ionescu, 2015; Mircea et al., 2015b). Automatic structure detection has strongly developed after 2019, likely due to the proliferation of image recognition algorithms based on convolutional neural networks (CNNs) and segmentation models such as Multi-Scale Region-Based CNNs, especially since 2021. The heterogeneity of models used in archaeology contrasts with other fields such as medical imaging applications, where CNNs and generative adversarial networks (GANs) are predominantly used (Barragán-Montero et al., 2021).
@@
Our work also highlighted some methodological concerns likely arising from the limited background of many archaeologists in machine learning, as well as of some programmers in relation to archaeological problems. Davis and Douglass have already pointed out this disconnect in remote sensing applications in archaeology, noting a “disconnect between remote sensing applications and anthropological theory(Davis and Douglass, 2020). Another critical aspect is the reliability and variability of the benchmarks themselves, which are often poor, leading to additional problems in evaluating the models (Tennie, 2023). Furthermore, the suitability of the chosen machine learning methodology to answering the scientific question sought after is crucial: in many cases, statistical methods are sufficient to provide an answer, possibly rendering machine learning methods superfluous. To address the challenges of applying these methods to archaeological questions, we have designed a flowchart of suggested best practices [Fig.˜11] to help archaeologists develop a coherent and effective approach. This small tool aims to bridge the gap between archaeological research and machine learning, promoting a more integrated and informed application of these technologies in the field.
@@
Regarding the future of machine learning in archaeology, three key phenomena emerge. Firstly, archaeologists’ ongoing training in digital methods will lead to a more comprehensive workflow and new applications in sub-fields of archaeology little represented in our current corpus (e.g., ethnoarchaeology, archaeobotany). Improving skills and knowledge will also reduce possible methodological issues, and provide a more direct interaction between the original research questions and the available avenues to solve them. Secondly, the constant development of new models will leave space for new proof-of-concept studies with more complex and wider datasets, but also allow for results that should improvement in results in some domains (cultural heritage, text analysis, automatic structure detection, and artefact classification). Thirdly, the recent boom in large language models and generative AI will change research in archaeology at different levels, as it does already for other fields of academia (Andersen et al., 2025). Either by allowing the creation of helpful tools for data curation, teaching tools, or multivariate analysis, these models are easy to use and have already been adopted by many in their daily life (though ethical issues remain; see Hagendorff 2024). One trend whose path remains uncertain is the possible formation of a reflexive community for AI and machine learning applications in archaeology (e.g. such as the COST action MAiA: MAIA 2025; or CAA machine learning special interest group), leading to more standardised practices and protocols with ethical guidelines.
@@
Despite the many concerns on applying machine learning in archaeology, these methods can be very powerful tools, able to massively reduce the time, cost, and (whether positive or not) labour to process, analyse, and predict even large-scale and complex data [e.g. Kochkov et al. 2024]. The existing reported successes of machine learning in other areas of science, and its already increasing number of applications in archaeology in recent years, all show it is cementing its place in the future of the field. We are likely at – if not already in a post-digital area (Gattiglia, 2025) – a turning point in archaeology, as well as in other disciplines, where a new approach – machine learning – has the possibility to permeate all levels of archaeological practice. Like any other statistical tool, machine learning itself is neither inherently good nor bad; its impact on the field of archaeology depends on how it is used (Huggett, 2017). We should be mindful of the thinking derived from post-processual archaeology in the late 1970s (Hodder, 1986), developed in response to processual archaeology (Phillips and Willey, 1953). The answer of this post-processual archaeology was to: “explain the past rather than describing it(Yoffee and Fowles, 2010). Following this perspective, machine learning applications in archaeology should never be limited to data processing or analysis but should be integrated into a broader archaeological reflection with precise questions, albeit knowingly and carefully, making sure its limitations are well known, well reported, and well addressed.
@@
Author contributions: M.B., J.D.O.F., C.T. and T.S. conceived the idea for the manuscript and initiated and designed the research. M.B. and J.D.O.F., collected data and selected the articles. M.B. and J.D.O.F., performed the statistical analysis and interpreted the results. M.B. and J.D.O.F. wrote the manuscript, with critical input from all co-authors. All authors reviewed the manuscript.
@@
Acknowledgment: We would like to thank J. Padarian for providing help on scraping strategy and Á. Tejero-Cantero for the generous help and patience afforded during the earlier stage of the project.
@@
Funding: This work has received funding from the Deutsche Forschungsgemeinschaft (DFG) Collaborative Research Center (CRC) 1070 “ResourceCultures”, grant agreement n°215859406 and from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program under grant agreement n°714658; (STONECULT project). We acknowledge support from the Open Access Publication Fund of the University of Tübingen.